# Binary Classification of Light and Dark Time Traces of a Transition Edge Sensor Using Convolutional Neural Networks

### v1

@@ V-A Comparison to Cut-Based Analysis
VI Background classification
We will use the CNN ensemble #2 from Section III-B to study the nature and origin of the remaining background set by the misclassified dark pulses (false positives). To do this, one needs to work with the $F_{1}$ score that quantifies how well the model classifies light pulses as light while avoiding misclassifying dark pulses and light. With the associated optimal thresholds, the ensemble #2 correctly classifies $97.5\%\pm 0.4\%$ of the light pulses (analysis efficiency) while misclassifying $2.00\%\pm 0.4\%$ of the dark pulses as light (Table II). The latter corresponds to the average of 157 misclassified dark pulses. It is likely that this subset of dark pulses is triggered by photonic events, making them difficult to distinguish from the light pulses. In fact, we have previously concluded that the limiting background source for our TES is fiber coupled near–1064 nm black-body photons which are indistinguishable from the light pulses given the energy resolution of the TES [20, 39, 40]. In order to confirm this, we begin by comparing the observed effective rate of the assumed near-1064 nm black-body background to theoretical predictions. Using the 1.77 d observation time derived in Section III-B, the observed background rate is $n_{\mathrm{b}}=157/1.77\,\mathrm{d}$ that equals approximately $1.02\,\mathrm{mHz}\pm 0.2\,\mathrm{mHz}$
Figure 8: (a) The average light pulse used for training and testing the CNN together with the average misclassified dark pulse. (b) PCA scatter plot showing the projection of feature vectors ($\tau_{\mathrm{rise}}$, $\tau_{\mathrm{decay}}$, $\chi_{\mathrm{Ph}}^{2}$, $\mathrm{Ph}_{\mathrm{FFT}}$, $\chi^{2}_{\mathrm{FFT}}$) onto the first two principal components (PC1 and PC2) for all of the light (true positives) and misclassified dark pulses (false positives).
@@ VI Background classification
$\dot{N}=\int d\Omega\int dA\int_{E_{1}}^{E_{2}}\frac{2}{h^{3}c^{2}}\cdot\frac{E^{2}}{\mathrm{e}^{E/kT}-1},$ (4)
where the first two integrals represent the solid angle ($\Omega$) and the area ($A$) over which the black-body radiation can enter the optical fiber, and $h$ is Planck’s constant, $c$ is the speed of light, $T$ is the temperature of the black-body radiation source and $k$ is Boltzmann’s constant. The integrals over $d\Omega$ and $dA$ are purely geometrical and can be estimated using the supplier provided specs of the used HI-1060 single mode fiber; numerical aperture $\mathrm{NA}=0.14$ and core radius $R=3.1\,\mu\mathrm{m}$. The solid angle is calculated as $\Omega=2\pi\cdot(1-\mathrm{cos}(\theta))$, where the acceptance angle of the fiber is $\theta=\mathrm{sin}^{-1}(\mathrm{NA})$. This results in $\Omega=0.062$. The corresponding area is simply $A=\pi R^{2}=3\cdot 10^{-11}\,\mathrm{m}^{2}$. The integration limits for the energy integral are set to $E\pm 3\sigma_{E}$ where the $E=1.165\,\mathrm{eV}$ corresponding to 1064 nm photons and $\sigma_{E}=0.088\,\mathrm{eV}$ based on the skewed Gaussian fit to the distribution of $\mathrm{Ph}_{\mathrm{FFT}}$ for light pulses. With these parameters, the integral in Eq. (4) at $T=295\,\mathrm{K}$ results in $\dot{N}=5.1\,\mathrm{mHz}$. The calculated rate is fivefold higher than what was observed by the CNN from the experimental data. However, the above presented calculation represents the theoretical maximum black-body rate and does not take into account various loss mechanisms present in experimental setup. In reality, this rate is lowered by the limited detection efficiency of the TES together with wavelength dependent transmission losses in the used mating sleeves and the fiber itself. In particular, fiber curling inside the cryostat, that was present in our experimental setup, has been observed to result in significant attenuation towards longer wavelength photons. The simulation of losses due to optical fiber, fiber curling and TES response in the same experimental setup used in this work has been recently addressed in Refs. [38, 40]. Using the associated simulation pipeline, we estimate a $0.57\,\mathrm{mHz}$ black-body background associated with cut-region $\mu_{\mathrm{m}}-3\sigma,\,\mu_{\mathrm{m}}+3\sigma]$. This corresponds better to the herein estimated value of $1.02\,\mathrm{mHz}\pm 0.2\,\mathrm{mHz}$. Considering the limitations of the simulation, the misclassified dark pulses seem indeed likely to result from near-1064 nm black-body photons coupled into the optical fiber based on the above presented rate comparison.
@@
The average misclassified dark pulse is illustrated in Fig. 8(a) together with the average light pulse. The shapes of the average pulses closely resemble each other suggesting a source of the same nature. In addition, Fig. 8(b) illustrates the PCA scatter plot showing the projections of the associated feature vectors ($\tau_{\mathrm{rise}}$, $\tau_{\mathrm{decay}}$, $\chi_{\mathrm{Ph}}^{2}$, $\mathrm{Ph}_{\mathrm{FFT}}$, $\chi^{2}_{\mathrm{FFT}}$) of the light and misclassified dark pulses onto the two main principal components. The majority of the misclassified dark pulses are clustered in the near vicinity of the light pulses. Only roughly 14 out of the 160 misclassified dark pulses are clearly outside the (red) cluster defined by light pulses. Thus, the vast majority of the misclassified dark pulses are most likely triggered by near-1064 nm photons originating from the black-body background. I.e., for the CNN they are indistinguishable from the light pulses given the energy resolution of the TES. Having these pulses distort the extrinsics (dark) dataset has detrimental effects on the training of the CNNs as the model is repeatedly being trained to classify an actual light pulse as dark. It is evident that this causes confusion in the training process of the CNN and limits the performance of the model. The presence of near-1064 nm black-body photon triggers in the dark dataset also explains why the CNNs are so susceptible towards the division of the dataset into training and testing sets (Fig. 3). How many of these, technically mislabeled dark pulses, end up in training or testing sets has a significant impact on both training and evaluation of the CNNs. This results in the observed high standard deviations in the evaluation metrics of the CNN ensemble #2 (Table II).
@@
VII Black-Body photons and training confusion
@@
In the previous section we concluded that the vast majority of the misclassified dark pulses are ultimately near-1064 nm photons. While these are physically indistinguishable from the light pulses given the energy resolution of the TES, training the CNNs to learn to classify these as dark pulses evidently causes confusion. The detrimental effects of this label noise have been widely studied [56, 57, 58]. While a small fraction of corrupted labels can result in improved performance by acting as a form of regularization, their effect on learning is generally detrimental [59]. This is particularly true for the herein studied CNNs, which are already regulated by the dropout layer.
@@
In consequence, the reported classifying performances of the CNNs in the previous sections should be considered as lower limits with room for further improvement when trained with an ideal, undistorted dataset. In order to demonstrate this, we relabel the target values of the 160 misclassified dark pulses from the previous section as light (target value $0\rightarrow 1$) and use this data for retraining an ensemble of 10 CNNs exactly as in Section VI (ensemble #2, see Fig. 5(b)). It should be noted that the 160 misclassified dark pulses form a subset of the testing data used in Section VI and consequently some near-1064 nm black-body photon triggers can still be left unidentified in the training set. Moreover, the 160 misclassified dark pulses had roughly 14 pulses which were clearly not clustered in the vicinity of the light pulses and thus these may not correspond to near-1064 nm black-body photons. Regardless, we relabel all of the 160 previously misclassified dark pulses as light, after which the vast majority of the near-1064 nm photon triggered dark counts should have been addressed. While the dataset still remains somewhat distorted, one expects to observe improvement in the CNN performance when trained with the relabeled dataset if there was training confusion present previously.
@@
Thus, we proceed in retraining the CNN ensemble #2 (Fig. 5(b)) using the dataset with 160 relabeled dark pulses based on Section VI. The overall dataset now contains $3898+160=4058$ light pulses and 8872160=87128872-160=8712 dark pulses which is then further divided into training and testing data according to Fig. (3). Upon doing this, the $F_{1}$ score improves from the previously estimated $\langle F_{1}\rangle=0.961\pm 0.005$ (Table II) to $\langle F_{1}\rangle=0.974\pm 0.004$. Evidently, the average $S$ score also improves due to additional background discrimination from $\langle S\rangle=0.95\pm 0.23$ to $\langle S\rangle=1.61\pm 0.58$ ($\epsilon_{\mathrm{a}}=85.5\%$ and $n_{\mathrm{b}}=0.17\,\mathrm{mHz}$), but still does not outperform the cut-based analysis after relabeling the presumed black-body triggers. The observed improvement in the $F_{1}$ score (see section III-B) is direct evidence that the performance of the CNN is limited by the training confusion caused by the presence of near-1064 nm black-body photon triggers in the measured extrinsic background.
VIII Conclusions and outlook
@@
We have aimed to improve the detection significance of a TES by analyzing the experimentally measured 1064 nm laser photon (light) and extrinsics background event (dark) triggered univariate time traces using a CNN based binary classifier. After extensive hyperparameter optimization, the CNN ensemble resulted in average detection significance of $\langle S\rangle=0.95\pm 0.23$, still being outperformed by our previously used (non-ML) cut-based analysis by 36%. While we have shown that this partly results from the training confusion introduced by the near-1064 nm black-body photons (indistinguishable from light) in the extrinsics background, we conclude that the binary classification based approach is not optimal analysis method for classifying the time traces. A CNN trained as a binary classifier does not seem to be able to learn how to suppress the noise of the time traces and is thus unable to improve the energy resolution of the TES, ultimately resulting in lower $S$ score when compared with the cut-based analysis.
In contrast to the binary classification scheme, we believe that regression based CNN would be better suitable for improving the detection significance and plan to investigate this further in future projects. The performance of the regression based models is presumably more sensitive to the underlying dataset, and thus the significance of dataset control and optimization should be stressed. Our recently published simulation framework for TES pulses provides particularly great opportunities for the required systematic dataset optimization [38, 40]. We also want to point out that the use of various ML models in unsupervised fashion, such as neural network based autoencoders, has shown great potential to address background suppression related tasks and should be further investigated [60].
@@
While there exist several potential post-data analysis methods that can improve the detection significance of the TES for the ALPS II experiment, we argue that reaching the black-body radiation limited [61, 40] ultra-low background of $10^{-5}\,\mathrm{Hz}$ ultimately requires the implementation of hardware-based background suppression methods. As already suggested in Ref. [61], the simplest way to do this is to apply a cryogenic narrow bandpass optical filter in front of the TES which effectively improves its energy resolution. We are currently building a cryogenic optical U-bench inside our dilution refrigerator enabling the implementation of such filter as a part of our TES setup.
IX Acknowledgements

### v2

@@ 5.1 Comparison to Cut-Based Analysis
6 Background classification
We will use the CNN ensemble #2 from Section 3.2 to study the nature and origin of the remaining background set by the misclassified dark pulses (false positives). To do this, one needs to work with the $F_{1}$ score that quantifies how well the model classifies light pulses as light while avoiding misclassifying dark pulses and light. With the associated optimal thresholds, the ensemble #2 correctly classifies $97.5\%\pm 0.4\%$ of the light pulses (analysis efficiency) while misclassifying $2.00\%\pm 0.4\%$ of the dark pulses as light (Table 2). The latter corresponds to the average of 157 misclassified dark pulses. It is likely that this subset of dark pulses is triggered by photonic events, making them difficult to distinguish from the light pulses. In fact, we have previously concluded that the limiting background source for our TES is fiber coupled near–1064 nm black-body photons which are indistinguishable from the light pulses given the energy resolution of the TES [Gimeno2023tes, Gimeno2024tes, Gimeno2025simulation]. In order to confirm this, we begin by comparing the observed effective rate of the assumed near-1064 nm black-body background to theoretical predictions. Using the 1.77 d observation time derived in Section 3.2, the observed background rate is $n_{\mathrm{b}}=157/1.77\,\mathrm{d}$ that equals approximately $1.02\,\mathrm{mHz}\pm 0.2\,\mathrm{mHz}$
Figure 9: (a) The average light pulse used for training and testing the CNN together with the average misclassified dark pulse. (b) PCA scatter plot showing the projection of feature vectors ($\tau_{\mathrm{rise}}$, $\tau_{\mathrm{decay}}$, $\chi_{\mathrm{Ph}}^{2}$, $V_{\mathrm{min,\,FFT}}$, $\chi^{2}_{\mathrm{FFT}}$) onto the first two principal components (PC1 and PC2) for all of the light (true positives) and misclassified dark pulses (false positives). The loading vectors associated with the principal components are $\mathbf{w}_{\mathrm{PC1}}=(2.9\cdot 10^{-4},\,1.0\cdot 10^{-3},\,0.99,-4.6\cdot 10^{-5},\,5.5\cdot 10^{-3})$ and $\mathbf{w}_{\mathrm{PC2}}=(-1.2\cdot 10^{-3},\,6.0\cdot 10^{-3},\,5.5\cdot 10^{-3},-1.4\cdot 10^{-2},\,0.99)$, again suggesting that the primary modes of variance are associated with the $\chi_{\mathrm{ph}}^{2}$ and $\chi_{\mathrm{FFT}}^{2}$ errors.
@@ 6 Background classification
$\dot{N}=\int d\Omega\int dA\int_{E_{1}}^{E_{2}}\frac{2}{h^{3}c^{2}}\cdot\frac{E^{2}}{\mathrm{e}^{E/kT}-1}\,dE,$ (4)
where the first two integrals represent the solid angle ($\Omega$) and the area ($A$) over which the black-body radiation can enter the optical fiber, and $h$ is Planck’s constant, $c$ is the speed of light, $T$ is the temperature of the black-body radiation source and $k$ is Boltzmann’s constant. The integrals over $d\Omega$ and $dA$ are purely geometrical and can be estimated using the supplier provided specs of the used HI-1060 single mode fiber; numerical aperture $\mathrm{NA}=0.14$ and core radius $R=3.1\,\mu\mathrm{m}$. The solid angle is calculated as $\Omega=2\pi\cdot(1-\mathrm{cos}(\theta))$, where the acceptance angle of the fiber is $\theta=\mathrm{sin}^{-1}(\mathrm{NA})$. This results in $\Omega=0.062$. The corresponding area is simply $A=\pi R^{2}=3\cdot 10^{-11}\,\mathrm{m}^{2}$. The integration limits for the energy integral are set to $E\pm 3\sigma_{E}$ where the $E=1.165\,\mathrm{eV}$ corresponding to 1064 nm photons and $\sigma_{E}=0.088\,\mathrm{eV}$ based on the skewed Gaussian fit to the distribution of $\mathrm{Ph}_{\mathrm{FFT}}$ for light pulses. With these parameters, the integral in Eq. (4) at $T=295\,\mathrm{K}$ results in $\dot{N}=5.1\,\mathrm{mHz}$. The calculated rate is fivefold higher than what was observed by the CNN from the experimental data. However, the above presented calculation represents the theoretical maximum black-body rate and does not take into account various loss mechanisms present in experimental setup. In reality, this rate is lowered by the limited detection efficiency of the TES together with wavelength dependent transmission losses in the used mating sleeves and the fiber itself. In particular, fiber curling inside the cryostat, that was present in our experimental setup, has been observed to result in significant attenuation towards longer wavelength photons. The simulation of losses due to optical fiber, fiber curling and TES response in the same experimental setup used in this work has been recently addressed in Refs. [RubieraGimeno2024optimizing, Gimeno2025simulation]. Using the associated simulation pipeline, we estimate a $0.57\,\mathrm{mHz}$ black-body background associated with cut-region $\mu_{\mathrm{m}}-3\sigma,\,\mu_{\mathrm{m}}+3\sigma]$. This corresponds better to the herein estimated value of $1.02\,\mathrm{mHz}\pm 0.2\,\mathrm{mHz}$. Considering the limitations of the simulation, the misclassified dark pulses seem indeed likely to result from near-1064 nm black-body photons coupled into the optical fiber based on the above presented rate comparison.
@@
The average misclassified dark pulse is illustrated in Fig. 9(a) together with the average light pulse. The shapes of the average pulses closely resemble each other suggesting a source of the same nature. In addition, Fig. 9(b) illustrates the PCA scatter plot showing the projections of the associated feature vectors ($\tau_{\mathrm{rise}}$, $\tau_{\mathrm{decay}}$, $\chi_{\mathrm{Ph}}^{2}$, $\mathrm{Ph}_{\mathrm{FFT}}$, $\chi^{2}_{\mathrm{FFT}}$) of the light and misclassified dark pulses onto the two main principal components. The majority of the misclassified dark pulses are clustered in the near vicinity of the light pulses. Only roughly 14 out of the 160 misclassified dark pulses are clearly outside the (red) cluster defined by light pulses. Thus, the vast majority of the misclassified dark pulses are most likely triggered by near-1064 nm photons originating from the black-body background. I.e., for the CNN they are indistinguishable from the light pulses given the energy resolution of the TES. Having these pulses distort the extrinsics (dark) dataset has detrimental effects on the training of the CNNs as the model is repeatedly being trained to classify an actual light pulse as dark. It is evident that this causes confusion in the training process of the CNN and limits the performance of the model. The presence of near-1064 nm black-body photon triggers in the dark dataset also explains why the CNNs are so susceptible towards the division of the dataset into training and testing sets (Fig. 4). How many of these, technically mislabeled dark pulses, end up in training or testing sets has a significant impact on both training and evaluation of the CNNs. This results in the observed high standard deviations in the evaluation metrics of the CNN ensemble #2 (Table 2).
@@
7 Black-body photons and training confusion
@@
In the previous section we concluded that the vast majority of the misclassified dark pulses are ultimately near-1064 nm photons. While these are physically indistinguishable from the light pulses given the energy resolution of the TES, training the CNNs to learn to classify these as dark pulses evidently causes confusion. The detrimental effects of this label noise have been widely studied [Liu2015classification, Song2022learning, Im2023binary]. While a small fraction of corrupted labels can result in improved performance by acting as a form of regularization, their effect on learning is generally detrimental [Lee2022binary]. This is particularly true for the herein studied CNNs, which are already regulated by the dropout layer.
@@
In consequence, the reported classifying performances of the CNNs in the previous sections should be considered as lower limits with room for further improvement when trained with an ideal, undistorted dataset. In order to demonstrate this, we relabel the target values of the 160 misclassified dark pulses from the previous section as light (target value $0\rightarrow 1$) and use this data for retraining an ensemble of 10 CNNs exactly as in Section 6 (ensemble #2, see Fig. 6(b)). It should be noted that the 160 misclassified dark pulses form a subset of the testing data used in Section 6 and consequently some near-1064 nm black-body photon triggers can still be left unidentified in the training set. Moreover, the 160 misclassified dark pulses had roughly 14 pulses which were clearly not clustered in the vicinity of the light pulses and thus these may not correspond to near-1064 nm black-body photons. Regardless, we relabel all of the 160 previously misclassified dark pulses as light, after which the vast majority of the near-1064 nm photon triggered dark counts should have been addressed. While the dataset still remains somewhat distorted, one expects to observe improvement in the CNN performance when trained with the relabeled dataset if there was training confusion present previously.
@@
Thus, we proceed in retraining the CNN ensemble #2 (Fig. 6(b)) using the dataset with 160 relabeled dark pulses based on Section 6. The overall dataset now contains $3898+160=4058$ light pulses and 8872160=87128872-160=8712 dark pulses which is then further divided into training and testing data according to Fig. (4). Upon doing this, the $F_{1}$ score improves from the previously estimated $\langle F_{1}\rangle=0.961\pm 0.005$ (Table 2) to $\langle F_{1}\rangle=0.974\pm 0.004$. Evidently, the average $S$ score also improves due to additional background discrimination from $\langle S\rangle=0.95\pm 0.23$ to $\langle S\rangle=1.61\pm 0.58$ ($\epsilon_{\mathrm{a}}=85.5\%$ and $n_{\mathrm{b}}=0.17\,\mathrm{mHz}$), but still does not outperform the cut-based analysis after relabeling the presumed black-body triggers. The observed improvement in the $F_{1}$ score (see section 3.2) is direct evidence that the performance of the CNN is limited by the training confusion caused by the presence of near-1064 nm black-body photon triggers in the measured extrinsic background. It should be noted that the presence of near-1064 nm black-body photon triggers in the dark dataset also imposes a physical limit on the classification performance of the cut-based analysis. Analogous to training of the CNNs, the skewed Gaussian distributions determining the cut regions were calculated using 1000 randomly selected light pulses, without any influence from the dark dataset containing the mislabeled dark pulses (see Section 5.1). After fixing the cut regions, the $S$ score was calculated using the rest of the data, including all of the near-1064 nm black-body photon triggers. Yet still, the cut-based analysis significantly outperforms the CNNs. As we previously concluded that inadequate hyperparameter optimization is not likely to limit the performance of the CNNs (see Section 3.3), this suggests that the CNN’s performance is limited by the training process due to the presence of mislabeled dark pulses.
8 Conclusions and outlook
@@
We have aimed to improve the detection significance of a TES by analyzing the experimentally measured 1064 nm laser photon (light) and extrinsics background event (dark) triggered univariate time traces using a CNN based binary classifier. After hyperparameter optimization, the CNN ensemble resulted in average detection significance of $\langle S\rangle=0.95\pm 0.23$, still being outperformed by our previously used (non-ML) cut-based analysis by 36%.
We have concluded that inadequate hyperparameter optimization is unlikely to limit the performance of the CNN. Meanwhile, our findings suggest that the limited performance can be attributed to training confusion introduced by the near-1064 nm black-body photon triggers in the extrinsics background. The CNN’s sensitivity to the used training data is particularly manifested as the large standard deviations in the calculated $S$ scores when compared with corresponding values obtained from cut-based analysis. Thus, the used experimental binary dataset seems to be inadequate for training the CNNs for improved noise suppression. Based on our results, we recommend further exploration of regression-based CNNs, with a strong focus on optimizing the size and structure of the training set. This includes, for example, systematically studying how the number and separation of distinct photon wavelengths affect the model’s regression performance when associating a pulse with a given wavelength. Our recently published simulation framework for TES pulses provides particularly great opportunities for the required systematic generation of datasets that correspond well to experimental data [RubieraGimeno2024optimizing, Gimeno2025simulation]. We also want to point out that the use of various ML models in unsupervised fashion, such as neural network based autoencoders, has shown great potential to address background suppression related tasks [Holl2019deep].
@@
While there exist several potential post-data analysis methods that can improve the detection significance of the TES for the ALPS II experiment, we argue that reaching the black-body radiation limited [Miller2007superconducting, Gimeno2025simulation] ultra-low background of $10^{-5}\,\mathrm{Hz}$ ultimately requires the implementation of hardware-based background suppression methods. As already suggested in Ref. [Miller2007superconducting], the simplest way to do this is to apply a cryogenic narrow bandpass optical filter in front of the TES which effectively improves its energy resolution. We are currently building a cryogenic optical U-bench inside our dilution refrigerator enabling the implementation of such filter as a part of our TES setup.
Data availibility
The datasets generated and/or analysed during the current study are available in the Dataset used for Binary Classification of Light and Dark Time Traces of a Transition Edge Sensor Using Convolutional Neural Networks repository, https://doi.org/10.5281/zenodo.17347454
Author contributions
E.R. implemented the machine learning models, made figures and wrote the manuscript. J.A.R.G. performed the cut-based analysis. E.R., G.O., J.A.R.G and C.S were involved with acquiring of experimental data. M.M. conceived the project together with K.-S.I, F.J and A.L, also providing supervision and support through the project. All authors discussed the results, provided feedback and approved the final version of the manuscript.
Declarations
Funding